OpenStudy (anonymous):

What is y=3x^2+1 in standard form?

OpenStudy (shadowlegendx):

Standard Form is Ax +By = C

OpenStudy (shadowlegendx):

You just have to use algebraic methods to get it into that format

OpenStudy (shadowlegendx):

Lets begin

OpenStudy (shadowlegendx):

y=3x^2+1 Look at how itʻs arranged compared to standard form y=3x^2+1 Ax + By = C All we have to do is line everything up, and abide by some rules, such as Ax canʻt being negative

OpenStudy (shadowlegendx):

y=3x^2+1 First we subtract 3x^2 from both sides -3x^2 + y = 1 Now we multiply both sides by -1 3x^2 - y = -1

OpenStudy (shadowlegendx):

See how it lines up with Ax + By = C 3x^2 - y = -1 Ax + By = C The signs can be different, except for Ax, where it canʻt be negative

OpenStudy (anonymous):

Thank you! This was the only problem I had that I couldn't figure out

OpenStudy (shadowlegendx):

No problem

OpenStudy (shadowlegendx):

Also, if you come across fractions in this, make sure to remove them using multiplication, as you canʻt have those suckas in there ;p

OpenStudy (shadowlegendx):

@ShadowLegendX me if you need help again

OpenStudy (anonymous):

Will do!